the points A, B and C represent the complex numbers
z1, z2, (1-i)z1 + iz2. respectively on complex plane. The triangle ABC is (right triangle, isoceles.........)

Grade:Upto college level

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

length of AB = |z1-z2|

length of BC = |1-i||z1-z2| = 2^0.5*|z1-z2|

length of AC = |-i||z1-z2| = |z1-z2|

this implies |AB| = |AC| and |BC| = square root of sum of squares of both AB and AC

so the triangle is right angle isosceles triangle

i hope it explains !!!